# Effective Whitney Stratification of Real Algebraic Varieties

### v3

@@ 1. Introduction
The algorithmic tools available for studying algebraic varieties differ substantially depending on whether one works over the real numbers or the complex numbers. For real varieties, the cylindrical algebraic decomposition algorithm [2, 7, 8, 15, 53] serves as a fundamental tool, while complex varieties are often studied through Gröbner basis techniques [11, 16, 41]. While these techniques are distinct both in approach and in the type of output produced, they share a worst-case complexity bound: both are doubly exponential in the number of variables. Nevertheless, a clear contrast exists between the complex and real cases in terms of the practical difficulty of solving fundamental problems; consider, for instance, the task of obtaining a closed form description of the connected components of a variety. This task is typically accomplished much more effectively over $\mathbb{C}$ than over $\mathbb{R}$ by computing a primary decomposition of the defining ideal via Gröbner basis calculation. Such disparities compel us to solve algorithmic tasks in real algebraic geometry by adapting computational methods which are available for complex varieties.
Real varieties arise more naturally in applications than their complex counterparts. Unfortunately, the geometry of a real variety is not readily revealed by its defining polynomials: for instance, every real variety is the vanishing locus of a single polynomial, whereas the minimal number of defining polynomials forms an important (and non-trivial) invariant for complex varieties. As such, one often seeks methods to subdivide a given variety into simpler regions which fit together coherently. And indeed, cylindrical algebraic decomposition provides precisely such a subdivision into contractible cells. One of our goals in this work is to outline what we view as a key step in an alternative framework for studying broad classes of applied and computational problems in real algebraic geometry.
@@ A Whitney stratification of a subset $X\subset\mathbb{R}^{n}$ is any locally-finite decomposition
into smooth, connected nonempty manifolds $M_{\alpha}\subset X$ called strata, so that every pair $(M_{\alpha},M_{\beta})$ satisfies Condition (B). It is known that every (real or complex) algebraic variety $X$ admits a Whitney stratification such that for each dimension $i$ the union $X_{i}$ of all strata of dimension $\leq i$ forms a subvariety (see e.g. [51, page 36–38]). One may therefore identify any such stratification of a $k$-dimensional variety $X$ with a flag
$X_{\bullet}=\big{(}\varnothing\subset X_{0}\subset X_{1}\subset\cdots\subset X_{k-1}\subset X_{k}=X\big{)}$
@@
One key consequence of imposing Condition (B) on strata pairs is equisingularity: a sufficiently small tubular neighborhood around a given stratum forms a (locally trivial) stratified fiber bundle over that stratum [39, Cor 10.6]. As a consequence, small neighborhoods in $X$ of two points lying in the same stratum $M_{\alpha}$ are stratified homeomorphic [^1] A homeomorphism $\phi:X\to Y$ of stratified spaces is said to be stratified whenever it maps each stratum $S\subset X$ diffeomorphically onto a stratum $\phi(S)\subset Y$. to each other. Consider, for instance, the (real) Whitney Umbrella from Figure 1. This is the variety $X\subset\mathbb{R}^{3}$ defined by $x^{2}-y^{2}z=0$, whose singular locus is the (positive) $Z$-axis (depicted as a purple line in the figure). If we take a small neighbourhood in $X$ of a point $(0,0,z)$ lying on the $Z$-axis, then its topology changes depending on whether $z$ is positive, negative or zero. Thus, the origin must constitute a stratum separate from the rest of the $Z$-axis in any valid Whitney stratification of $X$.
@@
The rest of this paper is organised as follows. In Section 2 we summarise relevant results and algorithms from [26] for stratifying complex varieties. The only new material here lies in Section 2.3, where we are able to remove a genericity hypothesis for map stratifications. Section 3 introduces a new WhitneyPolar algorithm for complex Whitney stratifications which is based on polar varieties. In Section 4 we show how both our old and new complex stratification algorithms may be used to stratify real varieties and algebraic morphisms between them. Our task in Section 5 is to similarly stratify basic closed semialgebraic sets whose inequality loci are open subsets of the ambient space — these often arise in practical contexts owing to boundedness or positivity constraints. Finally, in Section 6, we examine the complexity and experimental performance of WhitneyPolar.
@@ Acknowledgements
The material in this article is based upon work supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-22-1-0462, managed by Dr. Frederick Leve; both MH and VN were partially supported by this award and would like to gratefully acknowledge this support. Finally, we thank the two anonymous referees whose suggestions greatly improved our paper.
@@ 2. Whitney Stratifications from Conormal Ideals
In this section we begin with a brief review of the stratification algorithms for complex varieties and their morphisms. Section 2.1 and Section 2.2 are a review of the results of [26], restated in a form better suited to the current presentation. On the other hand, Section 2.3 is mostly new and makes a key generalization of the map stratification algorithm of [26, Section 6]; it allows us to drop the genericity requirement from [26, Definition 6.3], so that the Thom-Boardman flag of the given polynomial map $f:X\to Y$ need not intersect $X$ transversely [^3] See [19, Chapter VI] for a detailed treatment of Thom-Boardman flags.. Instead, we consider arbitrary polynomial maps $f:X\to Y$ between complex affine varieties $X\subset\mathbb{C}^{n}$ and $Y\subset\mathbb{C}^{m}$, and derive (in Lemma 2.5) an expression for the locus of smooth points of $X$ where the rank of $f$’s Jacobian drops below its generic value.
@@ 2.1. Stratification of Complex Varieties
Consider a pure $k$-dimensional complex variety, that is a variety all of whose irreducible components are of dimension $k$, $X=\mathbf{V}_{\mathbb{C}}(I_{X})\subset\mathbb{C}^{n}$ defined by a radical ideal $I_{X}=\langle f_{1},\dots,f_{r}\rangle$ in $\mathbb{C}[x_{1},\dots,x_{n}]$. Let $X_{\rm reg}$ denote the (open, dense) manifold of smooth points in $X$ and let $X_{\rm sing}=X-X_{\rm reg}$ be the singular locus of $X$; recall $X_{\rm sing}$ is a closed proper subvariety of $X$, see e.g. [48, Section 6.1] or  [40, page 26–27]. The conormal variety of $X$ is the subvariety of $X\times\mathbb{P}^{n-1}$ given by
@@
In other words, ${\bf Con}(X)$ is obtained by closing the set of pairs $(p,\xi)$ consisting of points $p\in X_{\rm reg}$ and $\xi\in\mathbb{P}^{n-1}$ such that the inner product $\left\langle{\xi,v}\right\rangle$ is zero for all $v\in T_{p}X_{\rm reg}$. We note the later coordinate is taken to be projecive as we identify all constant multiples of a particular unit direction vector defining the normal to a plane. The conormal variety comes endowed with a canonical projection map $\kappa_{X}:{\rm\bf Con}(X)\to X$ induced by the coordinate projection $\mathbb{C}^{n}\times\mathbb{P}^{n-1}\to\mathbb{C}^{n}$. Work in the ring $\mathbb{C}[x_{1},\dots,x_{n}][\xi_{1},\dots,\xi_{n}]$ and set
@@
The conormal variety ${\rm\bf Con}(X)=\mathbf{V}_{\mathbb{C}}(I_{{\rm\bf Con}(X)})$ is defined by the radical ideal obtained via the saturation $I_{{\rm\bf Con}(X)}=(I_{X}+K):J^{\infty}$ — here $K$ is the ideal generated by all $(n-k+1)\times(n-k+1)$ minors of the matrix $\mathscr{K}$, while $J$ is similarly defined by the $(n-k)\times(n-k)$ minors of the matrix $\mathscr{J}$. Note that the singular locus $X_{\rm sing}$ is precisely $\mathbf{V}_{\mathbb{C}}(J)$. A definition of ideal saturation and a discussion of how it is computed may be found in books such as [11, §4.4]. The following result [26, Theorem 3.1] forms the basis for efficient Whitney stratification of a given variety $X$. Recall that given an ideal $I$ its associated primes are the radicals of the ideals appearing in a minimal primary decomposition of $I$, see books such as [11, 40] for definitions and see [12] for a survey of methods for the computation of associated primes of an ideal via Gröbner basis.
@@ 8 Return $X_{\bullet}$
    Let $V_{\bullet}$ and $W_{\bullet}$ be nested sequences of subvarieties of a common variety, with the length $p$ of $V_{\bullet}$ larger than the length $q$ of $W_{\bullet}$. Then Merge$(V_{\bullet},W_{\bullet})$ creates a new sequence $U_{\bullet}$ of length $p$ via $U_{i}:=V_{i}\cup W_{\max(i,q)}$.
@@ 2.2. Flag-Subordinate Stratifications
Our main motivation for considering flag subordinate stratification is to develop algorithms to stratify algebraic maps between varieties. It is shown in [26, §5] that a flag subordinate stratification of a complex algebraic variety can always be computed using the WhitStratFlag algorithm described below. There are two main subroutines, the first of which is called InducedFlag. This takes as input a subvariety $W\subset X$ and a flag $\mathbf{F}_{\bullet}$ on $X$, and outputs the restriction of $\mathbf{F}_{\bullet}$ to $W$.
InducedFlag$(W,\mathbf{F}_{\bullet})$
@@ 5         If $\dim\mathbf{V}_{\mathbb{C}}(K)<\dim Y$
6               Merge $Y_{\bullet}$ with InducedFlag$(\mathbf{V}_{\mathbb{C}}(K),\mathbf{F}_{\bullet}X)$
@@ 7 Return $Y_{\bullet}$
Finally, here is the promised WhitStratFlag algorithm which makes use of these two subroutines. This takes as input a variety $X$ and a flag $\mathbf{F}_{\bullet}$ defined on $X$. The output $X_{\bullet}$ is guraranteed to be a Whitney stratification subordinate to $\mathbf{F}_{\bullet}$ in the sense of Definition 2.2 — see [26, Sec 5] for details.
WhitStratFlag$(X,\mathbf{F}_{\bullet})$
@@ 3 Set $X_{d}=X_{\rm sing}$ for all $d$ in ${\left\{{\mu,\mu+1,\ldots,k-1}\right\}}$
4 Merge $X_{\bullet}$ with InducedFlag$(X_{\rm sing},\mathbf{F}_{\bullet}X)$
@@ 2.3. Stratifying Complex Algebraic Morphisms
Maps of Whitney stratified spaces are typically required to satisfy additional criteria beyond smoothly sending strata to strata — see [6, Def 3.5.1] or [20, Part I, Ch 1.7] for instance.
@@
Let $\mathscr{X}_{\bullet}$ and $\mathscr{Y}_{\bullet}$ be Whitney stratifications of topological spaces $\mathscr{X}$ and $\mathscr{Y}$. A continuous function $\phi:\mathscr{X}\to\mathscr{Y}$ is stratified with respect to $\mathscr{X}_{\bullet}$ and $\mathscr{Y}_{\bullet}$ if for each stratum $M\subset\mathscr{X}$ there exists a a stratum $N\subset\mathscr{Y}$ satisfying two requirements:
@@
The second requirement of Definition 2.3 ensures the following crucial property via Thom’s first isotopy lemma [39, Prop 11.1]. If $\phi$ is a proper map – namely, if the inverse image of every compact subset of $Y$ is compact in $X$ – then for every stratum $N\subset Y$, the restriction of $\phi$ forms a locally trivial fiber bundle from $\phi^{-1}(N)$ to $N$.
Fix varieties $X\subset\mathbb{C}^{n}$ and $Y\subset\mathbb{C}^{m}$ along with a polynomial map $f:\mathbb{C}^{n}\to\mathbb{C}^{m}$ that sends points of $X$ to points of $Y$. In [26, Section 6], we described an effective algorithm for stratifying any $f:X\to Y$ whose Thom-Boardman flag [^5] See [19, Chapter VI]. intersects $X$ transversely. Here we will describe a new modification which circumvents this transversality requirement. The first step towards this goal is the following result.
@@
rank $\text{rank }\textbf{A}(x)\geq n-\dim X.$
@@
    rank $\text{rank }\text{\bf J}f^{*}(x)=0$,
@@
Similarly, rank $\text{rank }\text{\bf J}f^{*}(x)=i$ holds whenever rank $\text{rank }\mathbf{A}(x)=(n-\dim X)+i$; this rank condition on $\mathbf{A}(x)$ corresponds to the vanishing of all minors of size $(n-\dim X)+i+1$, as claimed above. ∎
@@
The first requirement of this definition forces $X_{\bullet}$ to depend on $\mathbf{S}_{\bullet}X$ whereas the second requirement makes $\mathbf{S}_{\bullet}X$ dependant on $X_{\bullet}$. We are therefore compelled to construct both $X_{\bullet}$ and $\mathbf{S}_{\bullet}X$ simultaneously via the algorithm described below. It will also be necessary in the sequel to produce WTB stratifications $(X_{\bullet},\mathbf{S}_{\bullet}X)$ for which $X_{\bullet}$ is subordinate to an auxiliary flag $\mathbf{F}_{\bullet}X$. Setting $\mathbf{F}_{i}X=X$ for all $i$, we obtain a WTB stratification as defined above. To simplify the presentation, we will denote this important special case as WTBStrat$(X,f)$.
WTBStratFlag$(X,f,\mathbf{F}_{\bullet}X)$

### v4

@@ 1. Introduction
Real varieties arise more naturally in applications than their complex counterparts. Unfortunately, the geometry of a real variety is not readily revealed by its defining polynomials: for instance, every real variety is the vanishing locus of a single polynomial, whereas the minimal number of defining polynomials forms an important (and non-trivial) invariant for complex varieties. As such, one often seeks methods to subdivide a given variety into simpler regions which fit together coherently. And indeed, cylindrical algebraic decomposition [2, 5, 6, 13, 47] provides precisely such a subdivision into contractible cells. One of our goals in this work is to outline what we view as a key step in an alternative framework for studying broad classes of applied and computational problems in real algebraic geometry.
@@ A Whitney stratification of a subset $X\subset\mathbb{R}^{n}$ is any locally-finite decomposition
into smooth, connected nonempty manifolds $M_{\alpha}\subset X$ called strata, so that every pair $(M_{\alpha},M_{\beta})$ satisfies Condition (B). It is known that every (real or complex) algebraic variety $X$ admits a Whitney stratification such that for each dimension $i$ the union $X_{i}$ of all strata of dimension $\leq i$ forms a subvariety (see e.g. [45, page 36–38]). One may therefore identify any such stratification of a $k$-dimensional variety $X$ with a flag
$X_{\bullet}=\big{(}\varnothing\subset X_{0}\subset X_{1}\subset\cdots\subset X_{k-1}\subset X_{k}=X\big{)}$ (1)
@@
One key consequence of imposing Condition (B) on strata pairs is equisingularity: a sufficiently small tubular neighborhood around a given stratum forms a (locally trivial) stratified fiber bundle over that stratum [36, Cor 10.6]. As a consequence, small neighborhoods in $X$ of two points lying in the same stratum $M_{\alpha}$ are stratified homeomorphic [^1] A homeomorphism $\phi:X\to Y$ of stratified spaces is said to be stratified whenever it maps each stratum $S\subset X$ diffeomorphically onto a stratum $\phi(S)\subset Y$. to each other. Consider, for instance, the (real) Whitney Umbrella from Figure 1. This is the variety $X\subset\mathbb{R}^{3}$ defined by $x^{2}-y^{2}z=0$, whose singular locus is the (positive) $Z$-axis (depicted as a purple line in the figure). If we take a small neighbourhood in $X$ of a point $(0,0,z)$ lying on the $Z$-axis, then its topology changes depending on whether $z$ is positive, negative or zero. Thus, the origin must constitute a stratum separate from the rest of the $Z$-axis in any valid Whitney stratification of $X$.
@@
The rest of this paper is organised as follows. In Section 2 we summarise relevant results and algorithms from [24] for stratifying complex varieties. The only new material here lies in Section 2.3, where we are able to remove a genericity hypothesis for map stratifications. Section 3 introduces a new WhitneyPolar algorithm for complex Whitney stratifications which is based on polar varieties. In Section 4 we show how both our old and new complex stratification algorithms may be used to stratify real varieties and algebraic morphisms between them. Our task in Section 5 is to similarly stratify basic closed semialgebraic sets whose inequality loci are open subsets of the ambient space — these often arise in practical contexts owing to boundedness or positivity constraints. Finally, in Section 6, we examine the experimental performance of WhitneyPolar.
@@ Acknowledgements
[…]
@@ 2. Whitney Stratifications from Conormal Ideals
In this section we begin with a brief review of the stratification algorithms for complex varieties and their morphisms. Section 2.1 and Section 2.2 are a review of the results of [24], restated in a form better suited to the current presentation. On the other hand, Section 2.3 is mostly new and makes a key generalization of the map stratification algorithm of [24, Section 6]; it allows us to drop the genericity requirement from [24, Definition 6.3], so that the Thom-Boardman flag of the given polynomial map $f:X\to Y$ need not intersect $X$ transversely [^3] See [16, Chapter VI] for a detailed treatment of Thom-Boardman flags.. Instead, we consider arbitrary polynomial maps $f:X\to Y$ between complex affine varieties $X\subset\mathbb{C}^{n}$ and $Y\subset\mathbb{C}^{m}$, and derive (in Lemma 2.5) an expression for the locus of smooth points of $X$ where the rank of $f$’s Jacobian drops below its generic value.
@@ 2.1. Stratification of Complex Varieties
Consider a pure $k$-dimensional complex variety, that is a variety all of whose irreducible components are of dimension $k$, $X=\mathbf{V}_{\mathbb{C}}(I_{X})\subset\mathbb{C}^{n}$ defined by a radical ideal $I_{X}=\langle f_{1},\dots,f_{r}\rangle$ in $\mathbb{C}[x_{1},\dots,x_{n}]$. Let $X_{\rm reg}$ denote the (open, dense) manifold of smooth points in $X$ and let $X_{\rm sing}=X-X_{\rm reg}$ be the singular locus of $X$; recall $X_{\rm sing}$ is a closed proper subvariety of $X$, see e.g. [43, Section 6.1] or  [37, page 26–27]. The conormal variety of $X$ is the subvariety of $X\times\mathbb{P}^{n-1}$ given by
@@
In other words, ${\bf Con}(X)$ is obtained by closing the set of pairs $(p,\xi)$ consisting of points $p\in X_{\rm reg}$ and $\xi\in\mathbb{P}^{n-1}$ such that the inner product $\left\langle{\xi,v}\right\rangle$ is zero for all $v\in T_{p}X_{\rm reg}$. We note the later coordinate is taken to be projective as we identify all constant multiples of a particular unit direction vector defining the normal to a plane. The conormal variety comes endowed with a canonical projection map $\kappa_{X}:{\rm\bf Con}(X)\to X$ induced by the coordinate projection $\mathbb{C}^{n}\times\mathbb{P}^{n-1}\to\mathbb{C}^{n}$. Work in the ring $\mathbb{C}[x_{1},\dots,x_{n}][\xi_{1},\dots,\xi_{n}]$ and set
@@
The conormal variety ${\rm\bf Con}(X)=\mathbf{V}_{\mathbb{C}}(I_{{\rm\bf Con}(X)})$ is defined by the radical ideal obtained via the saturation $I_{{\rm\bf Con}(X)}=(I_{X}+K):J^{\infty}$ — here $K$ is the ideal generated by all $(n-k+1)\times(n-k+1)$ minors of the matrix $\mathscr{K}$, while $J$ is similarly defined by the $(n-k)\times(n-k)$ minors of the matrix $\mathscr{J}$. Note that the singular locus $X_{\rm sing}$ is precisely $\mathbf{V}_{\mathbb{C}}(J)\cap X$. A definition of ideal saturation and a discussion of how it is computed may be found in books such as [9, §4.4]. The following result [24, Theorem 3.1] forms the basis for efficient Whitney stratification of a given variety $X$. Recall that given an ideal $I$ its associated primes are the radicals of the ideals appearing in a minimal primary decomposition of $I$, see books such as [9, 37] for definitions and see [10] for a survey of methods for the computation of associated primes of an ideal via Gröbner basis.
@@ 8 Return $X_{\bullet}$
[…]
@@ 2.2. Flag-Subordinate Stratifications
Our main motivation for considering flag subordinate stratification is to develop algorithms to stratify algebraic maps between varieties. It is shown in [24, §5] that a flag subordinate stratification of a complex algebraic variety can always be computed using the WhitStratFlag algorithm described below. There are two main subroutines, the first of which is called InducedFlag. This takes as input a subvariety $W\subset X$ and a flag $\mathbf{F}_{\bullet}$ on $X$, and outputs the restriction of $\mathbf{F}_{\bullet}$ to $W$.
[…]
@@ 5      If $\dim\mathbf{V}_{\mathbb{C}}(K)<\dim Y$
[…]
@@ 7 Return $Y_{\bullet}$
Finally, here is the promised WhitStratFlag algorithm which makes use of these two subroutines. This takes as input a variety $X$ and a flag $\mathbf{F}_{\bullet}$ defined on $X$. The output $X_{\bullet}$ is guraranteed to be a Whitney stratification subordinate to $\mathbf{F}_{\bullet}$ in the sense of Definition 2.2 — see [24, Sec 5] for details.
[…]
@@ 3 Set $X_{d}=X_{\rm sing}$ for all $d$ in $\left\{{\mu,\mu+1,\ldots,k-1}\right\}$}
[…]
@@ 2.3. Stratifying Complex Algebraic Morphisms
Maps of Whitney stratified spaces are typically required to satisfy additional criteria beyond smoothly sending strata to strata — see [4, Def 3.5.1] or [17, Part I, Ch 1.7] for instance.
@@
Let $\mathscr{X}_{\bullet}$ and $\mathscr{Y}_{\bullet}$ be Whitney stratifications of topological spaces $\mathscr{X}$ and $\mathscr{Y}$. A continuous function $\phi:\mathscr{X}\to\mathscr{Y}$ is stratified with respect to $\mathscr{X}_{\bullet}$ and $\mathscr{Y}_{\bullet}$ if for each stratum $M\subset\mathscr{X}$ there exists a stratum $N\subset\mathscr{Y}$ satisfying two requirements:
@@
The second requirement of Definition 2.3 ensures the following crucial property via Thom’s first isotopy lemma [36, Prop 11.1]. If $\phi$ is a proper map – namely, if the inverse image of every compact subset of $Y$ is compact in $X$ – then for every stratum $N\subset Y$, the restriction of $\phi$ forms a locally trivial fiber bundle from $\phi^{-1}(N)$ to $N$.
Fix varieties $X\subset\mathbb{C}^{n}$ and $Y\subset\mathbb{C}^{m}$ along with a polynomial map $f:\mathbb{C}^{n}\to\mathbb{C}^{m}$ that sends points of $X$ to points of $Y$. In [24, Section 6], we described an effective algorithm for stratifying any $f:X\to Y$ whose Thom-Boardman flag [^5] See [16, Chapter VI]. intersects $X$ transversely. Here we will describe a new modification which circumvents this transversality requirement. The first step towards this goal is the following result.
@@
$\text{rank }\textbf{A}(x)\geq n-\dim X.$
@@
    $\text{rank }\text{\bf J}f^{*}(x)=0$,
@@
Similarly, $\text{rank }\text{\bf J}f^{*}(x)=i$ holds whenever $\text{rank }\mathbf{A}(x)=(n-\dim X)+i$; this rank condition on $\mathbf{A}(x)$ corresponds to the vanishing of all minors of size $(n-\dim X)+i+1$, as claimed above. ∎
@@
The first requirement of this definition forces $X_{\bullet}$ to depend on $\mathbf{S}_{\bullet}X$ whereas the second requirement makes $\mathbf{S}_{\bullet}X$ dependant on $X_{\bullet}$. We are therefore compelled to construct both $X_{\bullet}$ and $\mathbf{S}_{\bullet}X$ simultaneously via the algorithm described below. It will also be necessary in the sequel to produce WTB stratifications $(X_{\bullet},\mathbf{S}_{\bullet}X)$ for which $X_{\bullet}$ is subordinate to an auxiliary flag $\mathbf{F}_{\bullet}X$. Setting $\mathbf{F}_{i}X=X$ for all $i$, we obtain a WTB stratification as defined above. To simplify the presentation, we will denote this important special case as WTBStrat(X,f)(X,f).
[…]